The hypotenuse of a right triangle is 15 cm and one of its legs is 11 cm . What is the measurement of the other leg ?
Answer by 10301650(32)

c^2=a^2+b^2 --> FORMULA
c (hypotenuse)= 15
a= 11
b= ?
-----: you need to find the b. now, you need to transpose the a to the other side to solve for the other leg of the triangle which is the b.
c^2-a^2=b^2
.
15^2-11^2=b^2
.
225-121=b^2
.
104=b^2
.
sqrt(104)= sqrt(b^2)
.
10.2= b (ANSWER)